The value of |avector cross icap|^2+|avector cross jcap|^2 + |avector cross kcap| is
a^2
2a^2
3a^2
None of these

Grade:7

1 Answers

Apoorva Arora IIT Roorkee
9 years ago
Let vector a be
\vec{a}=a_{1}i+a_{2}j+a_{3}k
So first let us see\vec{a}\times i=a_{3}j-a_{2}k
so,|\vec{a}\times i|^{2}=a_{3}^{2}+a_{2}^{2}.
In similar ways
|\vec{a}\times j|^{2}=a_{1}^{2}+a_{3}^{2}
and
|\vec{a}\times k|^{2}=a_{1}^{2}+a_{2}^{2}
Therefore,
|\vec{a}\times k|^{2}+|\vec{a}\times j|^{2}+|\vec{a}\times k|^{2}=2a_{1}^{2}+2a_{2}^{2}+2a_{3}^{2}=2a^{2}.